OpenStudy (anonymous):

integral of 6 cot t/5 dt with limit from 5pie/3 to 5pie/3

OpenStudy (satan ):

Are they all the same limits? \(\Large \frac{5\pi}{3}\) to \(\Large \frac{5\pi}{3}\)?

OpenStudy (satan ):

If it is, it would be zero, because you're not really moving any direction..